### v1

@@ 9 Conclusions
This study presented an extensive empirical evaluation of intrinsically interpretable machine learning models, addressing the lack of systematic guidance for practitioners in the field. By benchmarking sixteen diverse methods across more than two hundred real-world datasets, we moved beyond aggregate performance metrics to uncover how model behavior is influenced by structural data characteristics such as sample size, dimensionality, linearity, and class imbalance. Our analysis confirms that while modern interpretable approaches can achieve competitive predictive accuracy, model selection should be carefully tailored to the specific properties of the data at hand rather than relying on a one-size-fits-all solution.
In the regression domain, we observed a clear performance hierarchy where EBM consistently demonstrated superior predictive accuracy. However, the stratification analysis revealed distinct niches for other methods. Symbolic Regression and IGANN proved robust in non-linear regimes, whereas traditional linear and regularized models remained effective choices for high-dimensional or linear problems. The distinction was less pronounced in classification tasks, where performance tiers were narrower. Nevertheless, EBM retained a general advantage, while GOSDT offered a compelling trade-off, delivering significantly sparser and more interpretable models than standard Decision Trees at the cost of increased training time. Our investigation into out-of-sample generalization further highlighted critical vulnerabilities in specific methods, such as the degradation of k-NN in high-dimensional spaces under distributional shifts, contrasting with the robustness of Logistic Regression in data-scarce or high-dimensional scenarios.
A significant insight emerging from this study is the tangible cost associated with interpretability. We found that models offering higher expressiveness or provable sparsity, such as IGANN, Symbolic Regression, and GOSDT, require training times that are orders of magnitude higher than their simpler counterparts. This computational overhead presents a practical barrier for real-time applications or large-scale data streams.
In this work, we deliberately focused on predictive performance and computational efficiency, and did not attempt to quantify interpretability itself, reflecting the absence of a universally accepted and objective interpretability metric. Developing rigorous, model-agnostic frameworks for measuring interpretability, therefore, remains an important direction for future research.
@@ A.1 Linear models: Logistic, Linear, Lasso, GLM, GAM
Extensions of linear models, such as LASSO regression, introduce regularization to promote sparsity, improving interpretability by selecting a subset of relevant features through an L1 penalty (tibshirani1996regression). Generalized Linear Models extend linear models to accommodate non-normal target distributions (e.g., Poisson, Gamma) via appropriate link functions (mccullagh2019generalized). Generalized Additive Models (GAMs) further enhance flexibility by allowing non-linear transformations of individual features, such as splines, or interactions between feature pairs, while maintaining an additive structure that preserves interpretability (hastie1986generalized). Different formulations and implementations of GAMs have been proposed. Explainable Boosting Machines (EBMs) extend the concept of GAMs using gradient boosting to capture complex feature interactions while preserving interpretability through per-feature visualizations and monotonic constraints (nori2019interpretml).
@@ A.2 Tree-based models: Decision Tree, GOSDT
Tree-based models provide a rule-based representation of decision processes. Decision Trees recursively partition the feature space into homogeneous regions, producing hierarchical rules that are easy to visualize (breiman1984classification). However, their interpretability and predictive performance can be compromised by suboptimal or overly large trees (lipton2018mythos), as classical algorithms often rely on greedy splitting heuristics. Optimal Sparse Decision Trees (OSDT) introduced a practical, bound-driven search for provably optimal trees on binary variables, improving sparsity compared to greedy methods (hu2019optimal). Subsequent work generalized and scaled this approach to handle continuous features and imbalanced data (lin2020generalized). More recent methods apply reference ensembles from black-box models to guide the search, substantially reducing computation while producing sparse trees that remain interpretable (mctavish2022fast).
@@ A.3 Probabilistic models: Naive Bayes
Naive Bayes is a probabilistic approach grounded in Bayes’ theorem, assuming conditional independence among features (hastie2009elements). Despite its simplicity, it is effective in many real-world scenarios and provides interpretable probabilistic outputs. Feature contributions to predictions can be directly assessed through the model’s likelihood functions, offering insights into how individual variables influence the predicted probabilities.
A.4 Instance-based: KNN
Instance-based methods, such as k-Nearest Neighbors (k-NN), make predictions by comparing new instances to stored examples in the training data (cover1967knn). k-NN is often considered interpretable because predictions can be traced back to the most similar observed instances, providing an intuitive understanding of how the model makes decisions.
While some may argue that instance-based methods offer only local explanations, similar to post-hoc methods like LIME (ribeiro2016should), k-NN’s predictions are fully determined by the nearest training samples, without any additional approximation or model transformation. Consequently, the behavior of the model across the entire dataset can, in principle, be understood by examining the stored instances, and the influence of each feature is directly observable through the similarity metric.
@@ A.5 Symbolic models: Symbolic Regression
Symbolic Regression seeks to discover explicit mathematical expressions that describe the relationship between input features and the target variable (la2021contemporary). By generating human-readable formulas, Symbolic Regression offers a high degree of interpretability, allowing domain experts to understand and validate the underlying patterns captured by the model. Methods in this category balance predictive accuracy with simplicity, often guided by constraints on expression complexity.
Appendix B Hyperparameter Ranges
Table 3 reports the hyperparameter ranges used for model selection across all interpretable methods. Search spaces were chosen to reflect standard practice while enforcing constraints that favor parsimonious and interpretable model configurations.
@@ Appendix B Hyperparameter Ranges
Linear Regression None None
@@
Table 3: Hyperparameter search spaces for classification and regression models.

### v2

@@ 7 Conclusions
This study presented an extensive empirical evaluation of intrinsically interpretable machine learning models, addressing the lack of systematic guidance for practitioners in the field. By benchmarking sixteen diverse methods across more than two hundred real-world datasets, we moved beyond aggregate performance metrics to uncover how model behavior is influenced by meta-features such as sample size, dimensionality, linearity, and class imbalance. Our analysis confirms that while modern interpretable approaches can achieve competitive predictive accuracy, model selection should be carefully tailored to the specific properties of the data at hand rather than relying on a one-size-fits-all solution.
Our central finding is a profound dichotomy between regression and classification tasks. In regression, a clear and predictable performance hierarchy emerges, led by EBM and Symbolic Regression, and can be reliably anticipated from dataset meta-features such as target correlation, non-linearity, and local smoothness. In contrast, classification exhibits no stable or recoverable performance hierarchy. Model performances are heavily overlapping, and meta-features fail to reliably predict pairwise algorithm superiority. This suggests that classifier performance cannot be organized into a consistent, meta-feature-driven structure, and instead depends on complex, dataset-specific class geometry rather than broad structural properties.
Cutting across both tasks, we identified a pervasive interpretability tax: methods that optimize for structural parsimony or expressive non-linearity can demand up to two orders of magnitude more training time than simpler baselines, a cost practitioners must weigh explicitly against the transparency benefits.
In this work, we deliberately focused on predictive performance and computational efficiency, and did not attempt to quantify interpretability itself, reflecting the absence of a universally accepted and objective interpretability metric. Developing rigorous, model-agnostic frameworks for measuring interpretability, therefore, remains an important direction for future research. Additionally, to accommodate the structural constraints of certain models and standardize our out-of-sample shift simulations, we restricted classification tasks to binary problems. While multiclass tasks can conceptually be decomposed into binary problems, native multiclass extensions introduce distinct complexities that may alter the observed performance hierarchy and computational tax. Future benchmarking should evaluate natively multiclass environments to verify if the classification predictability gap persists.
Ultimately, these findings contribute to a deeper empirical understanding of interpretable modeling for tabular data. They provide researchers and practitioners with a data-driven foundation for algorithm selection, emphasizing that the pursuit of inherently interpretable representations must be carefully balanced against specific task dynamics, out-of-sample robustness requirements, and available computational budgets.
[…]
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper
Acknowledgement
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support.
@@ A.1 Linear models: Logistic, Linear, Lasso, GLM, GAM
Extensions of linear models, such as LASSO regression, introduce regularization to promote sparsity, improving interpretability by selecting a subset of relevant features through an L1 penalty [tibshirani1996regression]. Generalized Linear Models extend linear models to accommodate non-normal target distributions (e.g., Poisson, Gamma) via appropriate link functions [mccullagh2019generalized]. Generalized Additive Models (GAMs) further enhance flexibility by allowing non-linear transformations of individual features, such as splines, or interactions between feature pairs, while maintaining an additive structure that preserves interpretability [hastie1986generalized]. Different formulations and implementations of GAMs have been proposed. Explainable Boosting Machines (EBMs) extend the concept of GAMs using gradient boosting to capture complex feature interactions while preserving interpretability through per-feature visualizations and monotonic constraints [nori2019interpretml].
@@ A.2 Tree-based models: Decision Tree, GOSDT
Tree-based models provide a rule-based representation of decision processes. Decision Trees recursively partition the feature space into homogeneous regions, producing hierarchical rules that are easy to visualize [breiman1984classification]. However, their interpretability and predictive performance can be compromised by suboptimal or overly large trees [lipton2018mythos], as classical algorithms often rely on greedy splitting heuristics. Optimal Sparse Decision Trees (OSDT) introduced a practical, bound-driven search for provably optimal trees on binary variables, improving sparsity compared to greedy methods [hu2019optimal]. Subsequent work generalized and scaled this approach to handle continuous features and imbalanced data [lin2020generalized]. More recent methods apply reference ensembles from black-box models to guide the search, substantially reducing computation while producing sparse trees that remain interpretable [mctavish2022fast].
@@ A.3 Probabilistic models: Naive Bayes
Naive Bayes is a probabilistic approach grounded in Bayes’ theorem, assuming conditional independence among features [hastie2009elements]. Despite its simplicity, it is effective in many real-world scenarios and provides interpretable probabilistic outputs. Feature contributions to predictions can be directly assessed through the model’s likelihood functions, offering insights into how individual variables influence the predicted probabilities.
A.4 Instance-based models: KNN
Instance-based methods, such as k-Nearest Neighbors (k-NN), make predictions by comparing new instances to stored examples in the training data [cover1967knn]. k-NN is often considered interpretable because predictions can be traced back to the most similar observed instances, providing an intuitive understanding of how the model makes decisions.
While some may argue that instance-based methods offer only local explanations, similar to post-hoc methods like LIME [ribeiro2016should], k-NN’s predictions are fully determined by the nearest training samples, without any additional approximation or model transformation. Consequently, the behavior of the model across the entire dataset can, in principle, be understood by examining the stored instances, and the influence of each feature is directly observable through the similarity metric.
@@ A.5 Symbolic models: Symbolic Regression
Symbolic Regression seeks to discover explicit mathematical expressions that describe the relationship between input features and the target variable [la2021contemporary]. By generating human-readable formulas, Symbolic Regression offers a high degree of interpretability, allowing domain experts to understand and validate the underlying patterns captured by the model. Methods in this category balance predictive accuracy with simplicity, often guided by constraints on expression complexity.
Appendix B Taxonomy of Interpretability
While the term “interpretability” is ubiquitous in current literature, it remains an ill-defined concept, often treated as a monolithic property rather than a multifaceted and often subjective one. While we distinguished intrinsic interpretability from post-hoc explainability in Section 1, the category of inherently interpretable models itself is far from homogeneous. Following the frameworks established by lipton2018mythos and molnar2020interpretable, we categorize our candidate models based on three core dimensions of transparency: simulatability, decomposability, and algorithmic transparency.
Simulatability
A model is considered simulatable if a human can reasonably inspect the entire model at once and calculate its output for a given input. This property is strictly constrained by the cognitive load required to trace a decision path. In our experimental design, this is operationalized through structural constraints. For tree-based methods, we ensure simulatability for standard Decision Trees by limiting maximum depth, and explicitly test GOSDT, which optimize for sparsity to produce provably minimal representations. Similarly, Symbolic Regression targets simulatability by searching for parsimonious mathematical expressions that can be inspected directly.
Decomposability
Decomposability requires that the individual components of a model, such as input features, weights, or shape functions, can be analyzed independently. The linear models in our benchmark satisfy this by assigning scalar weights to features, where the magnitude and sign directly indicate contribution. We extend this category to include GAMs (specifically EBM and IGANN). While computationally complex, these models remain decomposable because they isolate feature effects into additive shape functions or interaction terms that can be visualized and understood independently of the rest of the model.
Algorithmic Transparency
Algorithmic transparency refers to models where the mathematical process from input to output is fully open and can be scrutinized. This category includes $k$-NN (and instance-based algorithms in general) and Naive Bayes. While these models do not always produce a compact formula, they are transparent because their decision-making process is grounded in clear mathematical axioms: $k$-NN predictions are fully determined by the most similar stored instances, and Naive Bayes relies on explicit conditional independence and likelihood functions.
Appendix C Missing Values
In the main text, we described the preprocessing pipeline used in our experiments (Section 3.2). A key step in this process was the exclusion of rows and columns containing missing values. While discarding data, much like arbitrary imputation, can potentially introduce bias, we provide a quantification of these missing values here to demonstrate their minimal impact. Figures 1(a) and 1(b) illustrate the percentage of missing cells across all datasets.
As observed, the vast majority of the datasets contain no missing values. Specifically, only two regression datasets and seven classification datasets exhibit any missing data. Furthermore, the proportion of missing cells is negligible (under 1%) for all regression datasets and for all but three classification datasets. Consequently, we are confident that the deletion of these missing values did not meaningfully alter our findings.
((a)) Regression
((b)) Classification
Figure C.1: Percentage of missing values in the benchmark datasets.
Appendix D Hyperparameter Ranges
Table D.1 reports the hyperparameter ranges used for model selection across all interpretable methods. Search spaces were chosen to reflect standard practice while enforcing constraints that favor parsimonious and interpretable model configurations.
@@ Appendix D Hyperparameter Ranges
Linear Regr. None None
@@
Table D.1: Hyperparameter search spaces for classification and regression models.
Appendix E Complete Stratification Results
Figures E.1, E.2, E.3, and E.4 present the stratified results across all evaluated meta-features. As discussed in the main text, we observe that while regression provides good stratification between the methods, this is not the case for classification.
Figure E.1: Stratified results for the regression task (in-sample), part 1.
Figure E.2: Stratified results for the regression task (in-sample), part 2.
Figure E.3: Stratified results for the classification task (in-sample), part 1.
Figure E.4: Stratified results for the classification task (in-sample), part 2.
Appendix F Ablation Study and Analysis for Symbolic Regression
F.1 Training Time
Throughout our comparative evaluation, model training was restricted to a maximum time budget of 300 seconds. While the execution times for the majority of the evaluated methods fell safely below this threshold, Symbolic Regression was a notable exception; its execution time was intentionally set to reach this maximum allowable limit to accommodate its extensive search of the functional space. This naturally raises the question of whether this strict time limit artificially bottlenecked the method’s capabilities, and if allocating a larger computational budget would have yielded improved predictive performance. This concern is particularly relevant for datasets with large sample sizes, where SR’s ranks were observed to degrade compared to other methods.
To investigate this, we randomly selected 10 datasets from the tercile characterized by the largest sample sizes ($N$). We then re-evaluated SR on these datasets with a substantially increased time budget of 900 seconds (three times the original limit).
Figure F.1 illustrates the difference in $R^{2}$ scores achieved under the two time constraints. We observe that the performance difference is negligible, leading us to conclude that the 300-second time budget was not a primary limiting factor for SR’s predictive accuracy in our main experiments.
Figure F.1: Predictive accuracy comparison of Symbolic Regression given an extended time budget.
F.2 Model Size
Having ruled out computational constraints as the root cause of Symbolic Regression’s relative performance degradation on large datasets, we investigate the role of the model selection process. SR algorithms, including the PySR implementation used in our study, typically generate a Pareto front of candidate equations that balance predictive accuracy against mathematical complexity. The final equation is chosen based on a scoring criterion that penalizes complexity to enforce parsimony.
As illustrated in Figure F.2, the structural complexity of the selected equations remains remarkably low, even as the dataset sample size ($N$) increases significantly. In large-sample regimes, there is generally sufficient statistical evidence to support more complex models capable of capturing finer, non-linear patterns without a high risk of overfitting. However, the strict parsimony penalty prevents the SR algorithm from selecting these more expressive formulations.
Figure F.2: Distribution of selected Symbolic Regression model complexities across varying dataset sample sizes.
Appendix G Algorithm Selection and Classification Performance Landscape
As discussed in Section 5.4, we operationalized dataset meta-features to predict both pairwise algorithm superiority and overall dataset difficulty. Figure G.1 visualizes the predictive outcomes for the classification task. The pairwise predictability matrix in Figure 1(a) reveals that the Random Forest meta-model struggles to forecast the performance difference between most pairs of classifiers, frequently yielding negative $R^{2}$ values. This confirms that meta-features provide little reliable guidance for a priori algorithm selection in the classification domain. When predicting inherent dataset difficulty, the meta-model achieves a comparatively low $R^{2}$ score of 0.28. Figure 1(b) indicates that the most influential meta-features for determining task hardness are the Error Rate of the Nearest Neighbor Classifier (N3), the Entropy of the 5 nearest neighbors (E5), and the Maximum Fisher’s Discriminant Ratio (F1). This emphasizes that classification difficulty is primarily driven by local class overlap and decision boundary complexity rather than broad structural data properties, similar to regression.
In the main text, we demonstrated that classification performance lacks a stable hierarchy when evaluated using the F1 score. To confirm that this observation is not merely an artifact of the discrete nature of the F1 metric, we replicated the landscape analysis using the Brier score, which is a continuous proper scoring rule. Figure G.2 presents the results of this extended analysis. The PCA biplot in Figure 2(a) projects the dataset meta-features onto the first two principal components, with datasets colored by the best performing algorithm according to the Brier score. Similar to the F1 score results, there is no discernible clustering of winning models within the meta-feature space. Furthermore, the hierarchical clustering of the method performance profiles across datasets, shown in Figure 2(b), fails to identify cohesive performance clusters. Specifically, for the Brier score clustermap, the overall Silhouette score was 0.22, which further decreased to 0.15 when the algorithm was forced to find a number of clusters equal to the number of methods. These findings validate that the highly interleaved and idiosyncratic nature of classifier performance remains consistent across different evaluation metrics. Finally, as reported in Figure G.3, predicting dataset difficulty using the Brier score achieves an $R^{2}$ of 0.34, closely mirroring the low predictability observed with the F1 score, and fully supporting these conclusions.
((a)) Pairwise
((b)) Feature Importance
Figure G.1: (a) $R^{2}$ scores of the meta-models predicting the $\Delta$ performance between pairs of models in classification. (b) Random Forest feature importances for predicting dataset difficulty in classification.
((a)) PCA biplot of the dataset meta-features (classification) projected onto the first two principal components. Individual datasets are color-coded by the winning algorithm based on the Brier score.
((b)) Hierarchical clustering of method performance profiles across classification datasets (Brier score).
Figure G.2: Landscape analysis for classification tasks evaluated using the Brier score.
((a)) $R^{2}$ scores of the meta-models predicting the $\Delta$ performance (Brier score) between pairs of models in classification.
((b)) Random Forest feature importances for predicting dataset difficulty (Brier score) in classification tasks.
Figure G.3: Predictability of pairwise performance deltas and dataset difficulty in classification tasks (Brier score).
Appendix H Regression Performance Landscape
For completeness and to provide a direct comparison to the classification landscape discussed in the main text, we report the corresponding landscape analyses for the regression tasks.
((a)) PCA biplot of the dataset meta-features (regression) projected onto the first two principal components. Individual datasets are color-coded by the winning algorithm.
((b)) Hierarchical clustering of method performance profiles across regression datasets.
Figure H.1: Landscape analysis for regression tasks, displaying the PCA biplot of meta-features alongside the hierarchical clustering of method performance.
As noted in our primary analysis, regression datasets show a clearer performance hierarchy and stronger stratification induced by the meta-features than classification datasets. To visualize the underlying structure of the regression benchmark, we first project the dataset meta-features into a reduced dimensional space using PCA. Figure 1(a) illustrates the resulting biplot, with datasets color-coded by the winning regression algorithm.
To further quantify the grouping of the performance profiles, we also apply the same hierarchical clustering approach to the algorithms’ performances across all regression datasets. Figure 1(b) presents the resulting dendrogram and clustermap.
This clustering yielded a notably higher overall silhouette score of 0.54. When the algorithm was constrained to find a number of clusters equal to the number of evaluated methods, the silhouette score remained relatively stable at 0.30 (compared to 0.05 for classification). These visualizations and metrics corroborate the findings presented in the main text: the successes and failures of regression methods present distinct, measurable clusters that are more structurally cohesive than those observed in the classification benchmark.